On August 3, Cinco Construction purchased special-purpose equipment at a cost of $7,304,400. The useful life of the equipment was estimated to be eight years, with an estimated residual value of $61,090.

A. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the straight-line depreciation method (half-year convention).

B. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the 200 percent declining-balance method (half-year convention) with a switch to straight-line when it will maximize depreciation expense.
Answer #1

1) Depreciation expense using Straight line method

Depreciation expense= (Original cost-Residual value)/Estimated useful life

= $(7304400-61090)/8= $905413.75

Year Depreciation expense
1 (905413.75/2)= $452707
2 905414
3 905414
4 905414
5 905414
6 905414
7 905414
8 905414
9 (905413.75/2)= $452707
Total $7243312

2) Depreciation rate= 100*2/8= 25%

Year Depreciation expense Ending Book Value
1 (7304400*25%*1/2)= $913050 (7304400-913050)= $6391350
2 (6391350*25%)= 1597838 (6391350-1597838)= 4793512
3 (4793512*25%)= 1198378 (4793512-1198378)= 3595134
4 (3595134*25%)= 898784 (3595134-898784)= 2696350
5 (2696350*25%)= 674088 (2696350-674088)= 2022262
6 (2022262*25%)= 505566 (2022262-505566)= 1516696
7 485202 (1516696-485202)= 1031494
8 485202 (1031494-485202)= 546292
9 485202 (546292-485202)= 61090
Total $7243310

Depreciation under straight line method= $(1516696-61090)/3= $485202